### v2

# SMARTe: Slot-based Method for Accountable Relational Triple extraction

Abstract
Relational Triple Extraction (RTE) is a fundamental task in Natural Language Processing (NLP). However, prior research has primarily focused on optimizing model performance, with limited efforts to understand the internal mechanisms driving these models. Many existing methods rely on complex preprocessing to induce specific interactions, often resulting in opaque systems that may not fully align with their theoretical foundations. To address these limitations, we propose SMARTe: a Slot-based Method for Accountable Relational Triple extraction. SMARTe introduces intrinsic interpretability through a slot attention mechanism and frames the task as a set prediction problem. Slot attention consolidates relevant information into distinct slots, ensuring all predictions can be explicitly traced to learned slot representations and the tokens contributing to each predicted relational triple. While emphasizing interpretability, SMARTe achieves performance comparable to state-of-the-art models. Evaluations on the NYT and WebNLG datasets demonstrate that adding interpretability does not compromise performance. Furthermore, we conducted qualitative assessments to showcase the explanations provided by SMARTe, using attention heatmaps that map to their respective tokens. We conclude with a discussion of our findings and propose directions for future research.
SMARTe: Slot-based Method for Accountable Relational Triple extraction
Xue Wen Tan1 Stanley Kok2 1Asian Institute of Digital Finance, National University of Singapore 2School of Computing, National University of Singapore xuewen@u.nus.edu, skok@comp.nus.edu.sg
1 Introduction
Relational Triple Extraction (RTE) is a well-established and widely studied task in Natural Language Processing (NLP). Its primary objective is to automatically extract structured information such as names, dates, and relationships from unstructured text, thereby enhancing data organization and accessibility Nayak et al. (2021). These extracted relationships are represented as relational triples consisting of (Subject, Relation, Object).Such structured representations underpin a range of downstream applications, including knowledge graph construction, question answering, and information retrieval. RTE is often confused with Joint Entity and Relation Extraction (JERE) Zhang et al. (2017); Gupta et al. (2016); Miwa and Sasaki (2014), or the terms are used interchangeably by some scholars Sui et al. (2023); Li et al. (2021). However, there is a key distinction: JERE addresses both entity identification and relation extraction (i.e., ACE04, ACE05 and CoNLL04 datasets) Roth and Yih (2004), which require annotation of all entities in a text, regardless whether they participate in a relationship. In contrast, RTE focuses exclusively on entities that are part of a relationship, specifically a set of annotated relational triples, hence the two tasks are fundamentally different. In this paper, we focus solely on RTE.
To motivate our work, lets take a look at an example seen in Figure 1. Most approaches rely on fine-tuned BERT embeddings to form pair representations by permuting all token pairs and classifying their relationships. However, this methodology makes it difficult to trace which specific contextual information each pair has absorbed, for instance, whether "Barack" or "America" has incorporated the "president" context. Despite achieving superior F1 performance, these pair-based approaches lack interpretability, obscuring how models arrive at their predictions. In contrast, our approach uses triplet-based representation, which absorbs all relevant tokens to directly form relational triplets. This enables a more direct and interpretable encoding of relationships. We observed that most existing methods have excellent performance but with limited interpretability, which highlights a significant gap in current research. In this study, we introduce SMARTe111We will release our implementation after acceptance.: a Slot-based Method for Accountable Relational Triple Extraction, a transparent architecture designed to address interpretability in RTE. SMARTe incorporates intrinsic interpretability through the use of a slot attention mechanism. This approach ensures that every prediction from SMARTe can be explicitly traced back to its learned slot representations and the specific tokens contributing to each predicted relational triple as shown in Appendix A. This results in informed and transparent outputs, embodying the Accountability, which directly equates to explainability, in our model. Since the learned slots do not follow any inherent order, the RTE task is framed as a set prediction problem.
Although we illustrate our approach in the context of RTE, it is broadly applicable to other NLP tasks characterized by set-based target structures. Interpretability becomes especially valuable in high-stakes scenarios where transparency and trust are paramount Danilevsky et al. (2020). As stakeholders increasingly demand clear and interpretable explanations for AI-driven decisions, our approach effectively addresses these critical requirements. We introduce SMARTe, a Slot-based Method for Accountable Relational Triple extraction. Our contributions are as follows:
To the best of our knowledge, the proposed SMARTe framework is the first to introduce interpretability to relational triple extraction tasks by adapting the slot attention mechanism, originally developed for unsupervised learning in computer vision.
We conduct extensive experiments on two widely used datasets and demonstrate that our model achieves performance comparable to current state-of-the-art systems, all while offering interpretability.
We provide a qualitative assessment and demonstrate how slot attention facilitates explanations which allow users to understand the model’s reasoning behind its predictions.
We report zero-shot performance on RTE tasks using recent LLMs for future benchmarks.
2 Related Work
Recent advancements in relation triplet extraction (RTE) have investigated various architectures to better capture interactions between entities and relations. We group these approaches into three main categories: sequence-to-sequence (seq2seq) Methods, Tagging-Based Methods, and Pairwise-Based Methods.
Seq2seq methods treat triples as token sequences, leveraging an encoder-decoder framework akin to machine translation. CopyRE Zeng et al. (2018) uses a copy mechanism to generate relations and entities but struggles with multi-token entities. CopyMTL Zeng et al. (2020) addresses this by employing a multi-task learning framework. CGT Ye et al. (2021) introduces a generative transformer with contrastive learning to enhance long-term dependency and faithfulness. R-BPtrNet Chen et al. (2021) uses a binary pointer network to extract explicit and implicit triples, while SPN Sui et al. (2023) reframes relational triple extraction as a set prediction problem, utilizing a non-autoregressive decoder with iterative refinement for improved contextual representation.
Tagging-Based Methods, also known as sequence labeling methods, utilize binary tagging sequences to identify the start and end positions of entities and, in some cases, to determine relations. Early approaches, such as NovelTagging Zheng et al. (2017), introduced a tagging-based framework that reformulates joint extraction as a tagging problem, enabling direct extraction of entities and their relations. CasRel Wei et al. (2019) improves on this by first identifying all potential head entities and then applying relation-specific sequence taggers to detect corresponding tail entities. Recent methods, however, are not entirely sequence-based. For instance, BiRTE employs tagging only for entities. It employs a pipeline strategy in which entities are tagged as either subjects or objects, after which all possible subject-object pairs are generated and classified using a biaffine scorer. Similarly, PRGC Zheng et al. (2021) includes a component to predict potential relations, constraining subsequent entity recognition to the predicted relation subset. PRGC’s strong performance stems from its use of a global correspondence table, which effectively captures interactions between token pairs.
Pairwise-Based Methods focus on enhancing token-pair interaction representations to improve relation classification. These methods eliminate the need to explicitly predict head or tail entities, as token pairs classified as NA (no relationship) are directly discarded. Notably, approaches in this category have achieved state-of-the-art (SOTA) F1 performance. Early works often framed these methods as table-filling approaches. For example, GraphRel Fu et al. (2019) models entity-relation interactions through a relation-weighted Graph Convolutional Network. TPLinker Wang et al. (2020) redefines triple extraction as a token-pair linking task, utilizing a relation-specific handshaking tagging scheme to align boundary tokens of entity pairs. Similarly, PFN Yan et al. (2021) introduces a partition filter network that integrates task-specific feature generation to simultaneously model entity recognition and relation classification. Modern SOTA techniques further refine these approaches. UniRel Tang et al. (2022) employs a unified interaction map to effectively capture token-relation interactions and also incorporating relation-specific token information into the prediction process. Similarly, DirectRel Shang et al. (2022) reformulates relational triple extraction (RTE) as a bipartite graph linking task, focusing on generating head-tail entity pairs.
3 Approach
The architecture of our SMARTe model is depicted in Figure 2, consisting of three primary components: an encoder, a slot attention module and a structured relational triple extractor.
3.1 Encoder
The encoder transforms the input text into dense, contextual representations that capture the semantic and syntactic information necessary for downstream processing. In our implementation, we utilize pre-trained transformer-based models BERT-Base-Cased Devlin et al. (2019). The process begins with tokenizing the input text into subword units, which are then fed into the transformer. The output is a sequence of contextualized embeddings, each corresponding to a token in the input text. Given an input text sequence , where represents the -th token, the encoder transforms this sequence into a sequence of contextualized embeddings:
Here, , where is the length of a sequence of contextualized embeddings (including [CLS] and [SEP], two special start and end markers), with each corresponding to a token in the input text, and is the embedding size produced by the encoder for each token.
3.2 Slot Attention Module
Slot Attention is a neural network module designed for object-centric representation learning Locatello et al. (2020), enabling a model to decompose a scene into distinct entities or objects. It uses iterative attention mechanisms to map input data (like image features) to a fixed number of slots, which are learnable feature vectors representing objects or parts of the input. This approach is highly effective in unsupervised learning settings, as it disentangles objects without requiring explicit annotations. The general form of Slot Attention is as follow:
At iteration , the current slots are transformed into query vectors using a learnable weight matrix , while the input features , derived from the sequence of contextualized embeddings from the encoder , are projected into key and value vectors using and , respectively. Attention weights are computed as the dot product between and , followed by softmax to produce a probability distribution that ensures each slot attends to specific parts of the input. The slots are updated using a Gated Recurrent Unit (GRU) Chung et al. (2014), combining the previous slot representation with the weighted aggregation of values . This iterative process refines the slots, allowing the model to disentangle objects or entities in the input data. After completing the iterative process, the output is a refined slots , where is the number of slots and is the dimensionality of each slot.
However, the softmax function in slot attention can be too restrictive for relational triple extraction tasks, particularly when certain tokens, such as “Barack Obama,” are involved in multiple triples and need to be associated with multiple slots. For instance, in the triples (Barack Obama, Born In, Hawaii), (Barack Obama, President Of, United States), and (Barack Obama, Married To, Michelle Obama), the token Barack Obama serves as the subject in each case. Softmax enforces a normalization constraint where attention scores must sum to 1, causing the token’s contribution to be distributed across slots. This can dilute its impact and make it difficult for the model to maintain consistent associations across triples. To address this limitation, our SMARTe model employ the optimal transport variant outlined in Zhang et al. (2023), which is more relaxed and provides a flexible framework for assigning tokens to slots while preserving their relevance across multiple contexts. Specifically, this involves replacing equation 3 with the optimal transport algorithm Villani et al. (2009), as follows:
In equation 5, the initial transport cost is computed as the cosine similarity between and . Next, in equation 6, selects, from the set , the candidate matrix that minimizes the entropy of the initial cost function, . Here, iteratively normalizes the row and column sums to produce a doubly stochastic matrix. corresponds to Minimizing the Entropy of Sinkhorn. Finally, in equation 7, the chosen matrix is passed once more through the Sinkhorn operator to yield the final attention matrix . This “transport plan” preserves row- and column-stochastic constraints, providing a refined set of attention weights. For more details, we refer readers to the work of Zhang et al. (2023). We have also provided the results and analysis for the softmax variant for reference.
3.3 Structured Relational Triple Extractor
In the original Slot Attention paper Locatello et al. (2020), the approach was designed for fixed-size images. However, when applied to text, variable input lengths make direct coordinate prediction more challenging. To address this, we reformulate coordinate prediction as a one-hot sequence labeling task to identify the start and end positions. We perform matrix multiplication between the slot attention outputs and the sequence tokens to encode information from the slot attention representation into each token. A feedforward neural network with linear layers predicts the indices for entity head-tail pairs, along with their corresponding relationships, in a unified manner. Specifically, it identifies five key components: [subject-start (ss), subject-end (se), object-start (os), object-end (oe), and relationship (rs)], which collectively form a structured relational triple:
refers to the prediction for each component at the slot, with representing the slot embedding; the matrices , , where represents the total number of relation types in the RTE task and are learnable parameters. The function refers to the softmax operation and represents the sequence of contextualized embeddings.
Since the slots inherently lack ordering, directly comparing predictions with the ground truth during model training presents a challenge. To overcome this, the predictions must first be aligned with the ground truth. This alignment is achieved using the Hungarian matching algorithm Kuhn (1955). After optimally matching the predictions to the ground truth, cross-entropy loss is applied to each component to train the model:
where represents the optimal assignment that minimizes the total pairwise matching cost and denotes the prediction, represents the corresponding ground truth target at the slot and the indicator function ensuring only valid relational triples contribute to the loss.
4 Experimental Setup
We evaluate our model using two widely recognized benchmark datasets frequently employed in the existing literature on relational triple extraction: the New York Times (NYT) dataset Riedel et al. (2010) and the Web Natural Language Generation (WebNLG) dataset Gardent et al. (2017) (Please note that these are the only two benchmark datasets specifically designed for this task). Both datasets include two variants of annotation: Partial Matching and Exact Matching. In Partial Matching, only the head words of the ground truth entities are annotated, whereas in Exact Matching, the entire span of the entities is annotated. Therefore, Exact Matching offers a more precise and comprehensive representation of the task. It is worth mentioning that most prior studies primarily report results based on Partial Matching, often neglecting Exact Matching. In our study, we present results for both matching strategies to facilitate comparisons in future research. Detailed statistics for these datasets are provided in Table 5 and 6. Please refer to Appendix C for a detailed overview of the key hyperparameters and hardware used in our experiments.
5 Experimental Results
| Model | NYT* | NYT | ||||
| Prec. | Rec. | F1 | Prec. | Rec. | F1 | |
| CasRel Wei et al. (2019) | 89.7 | 89.5 | 89.6 | 90.1 | 88.5 | 89.3 |
| TPLinker Wang et al. (2020) | 91.3 | 92.5 | 91.9 | 91.4 | 92.6 | 92.0 |
| CGT Ye et al. (2021) | 94.7 | 84.2 | 89.1 | - | - | - |
| PRGC Zheng et al. (2021) | 93.3 | 91.9 | 92.6 | 93.5 | 91.9 | 92.7 |
| R-BPtrNet Chen et al. (2021) | 92.7 | 92.5 | 92.6 | - | - | - |
| BiRTE Ren et al. (2022) | 92.2 | 93.8 | 93.0 | 91.9 | 93.7 | 92.8 |
| DirectRel Shang et al. (2022) | 93.7 | 92.8 | 93.2 | 93.6 | 92.2 | 92.9 |
| UniRel Tang et al. (2022) | 93.5 | 94.0 | 93.7 | - | - | - |
| SPN Sui et al. (2023) | 93.3 | 91.7 | 92.5 | 92.5 | 92.2 | 92.3 |
| Phi4-14B‡ Abdin et al. (2024) | - | - | - | 3.3 | 6.1 | 4.0 |
| Gemma3-27B‡ Gemma-Team (2025) | - | - | - | 6.0 | 12.0 | 7.3 |
| Qwen3-32B‡ Qwen-Team (2025) | - | - | - | 3.7 | 8.2 | 4.7 |
| Llama3.3-70B‡ Grattafiori et al. (2024) | - | - | - | 7.0 | 15.2 | 8.9 |
| GPT 4o-Mini‡ OpenAI et al. (2024) | - | - | - | 3.3 | 6.3 | 3.9 |
| GPT 4.1-Mini‡ OpenAI et al. (2024) | - | - | - | 7.1 | 11.3 | 7.9 |
| SMARTe (Softmax) | 92.2 | 91.2 | 91.7 | 92.1 | 91.7 | 91.9 |
| SMARTe (Opt Transport) | 92.4 | 92.9 | 92.6 | 92.5 | 93.0 | 92.7 |
| SMARTe† (Opt Transport) | 92.5 | 93.3 | 92.9 | 92.7 | 93.1 | 92.9 |
| Model | WebNLG* | WebNLG | ||||
| Prec. | Rec. | F1 | Prec. | Rec. | F1 | |
| CasRel Wei et al. (2019) | 93.4 | 90.1 | 91.8 | - | - | - |
| TPLinker Wang et al. (2020) | 91.7 | 92.0 | 91.9 | 88.9 | 84.5 | 86.7 |
| CGT Ye et al. (2021) | 92.9 | 75.6 | 83.4 | - | - | - |
| PRGC Zheng et al. (2021) | 94.0 | 92.1 | 93.0 | 89.9 | 87.2 | 88.5 |
| R-BPtrNet Chen et al. (2021) | 93.7 | 92.8 | 93.3 | - | - | - |
| BiRTE Ren et al. (2022) | 93.2 | 94.0 | 93.6 | 89.0 | 89.5 | 89.3 |
| DirectRel Shang et al. (2022) | 94.1 | 94.1 | 94.1 | 91.0 | 89.0 | 90.0 |
| UniRel Tang et al. (2022) | 94.8 | 94.6 | 94.7 | - | - | - |
| SPN Sui et al. (2023) | 93.1 | 93.6 | 93.4 | - | - | - |
| Phi4-14B‡ Abdin et al. (2024) | - | - | - | 13.5 | 16.4 | 14.3 |
| Gemma3-27B‡ Gemma-Team (2025) | - | - | - | 24.6 | 31.4 | 26.6 |
| Qwen3-32B‡ Qwen-Team (2025) | - | - | - | 15.7 | 19.2 | 16.8 |
| Llama3.3-70B‡ Grattafiori et al. (2024) | - | - | - | 25.7 | 31.3 | 27.3 |
| GPT 4o-Mini‡ OpenAI et al. (2024) | - | - | - | 24.0 | 28.8 | 25.3 |
| GPT 4.1-Mini‡ OpenAI et al. (2024) | - | - | - | 29.6 | 32.8 | 30.1 |
| SMARTe (Softmax) | 92.8 | 92.5 | 92.7 | 85.2 | 83.8 | 84.5 |
| SMARTe (Opt Transport) | 93.4 | 93.4 | 93.4 | 86.0 | 85.6 | 85.8 |
| SMARTe† (Opt Transport) | 93.7 | 94.0 | 93.9 | 86.6 | 86.0 | 86.3 |
| Model | Normal | EPO | SEO | N=1 | N=2 | N=3 | N=4 | N5 |
| CasRel Wei et al. (2019) | 87.3 | 92.0 | 91.4 | 88.2 | 90.3 | 91.9 | 94.2 | 83.7 |
| TPLinker Wang et al. (2020) | 90.1 | 94.0 | 93.4 | 90.0 | 92.8 | 93.1 | 96.1 | 90.0 |
| PRGC Zheng et al. (2021) | 91.0 | 94.5 | 94.0 | 91.1 | 93.0 | 93.5 | 95.5 | 93.0 |
| BiRTE Ren et al. (2022) | 91.4 | 94.2 | 94.7 | 91.5 | 93.7 | 93.9 | 95.8 | 92.1 |
| DirectRel Shang et al. (2022) | 91.7 | 94.8 | 94.6 | 91.7 | 94.1 | 93.5 | 96.3 | 92.7 |
| UniRel Tang et al. (2022) | 91.6 | 95.2 | 95.3 | 91.5 | 94.3 | 94.5 | 96.6 | 94.2 |
| SPN Sui et al. (2023) | 90.8 | 94.1 | 94.0 | 90.9 | 93.4 | 94.2 | 95.5 | 90.6 |
| SMARTe (Softmax) | 90.3 | 93.4 | 92.9 | 90.3 | 93.1 | 93.3 | 95.5 | 85.3 |
| SMARTe (Opt Transport) | 90.7 | 94.5 | 94.3 | 90.6 | 93.6 | 94.4 | 96.2 | 90.7 |
| SMARTe† (Opt Transport) | 90.7 | 95.1 | 94.7 | 90.7 | 94.2 | 94.3 | 96.5 | 91.4 |
| Model | Normal | EPO | SEO | N=1 | N=2 | N=3 | N=4 | N5 |
| CasRel Wei et al. (2019) | 89.4 | 94.7 | 92.2 | 89.3 | 90.8 | 94.2 | 92.4 | 90.9 |
| TPLinker Wang et al. (2020) | 87.9 | 95.3 | 92.5 | 88.0 | 90.1 | 94.6 | 93.3 | 91.6 |
| PRGC Zheng et al. (2021) | 90.4 | 95.9 | 93.6 | 89.9 | 91.6 | 95.0 | 94.8 | 92.8 |
| BiRTE Ren et al. (2022) | 90.1 | 94.3 | 95.9 | 90.2 | 92.9 | 95.7 | 94.6 | 92.0 |
| DirectRel Shang et al. (2022) | 92.0 | 97.1 | 94.5 | 91.6 | 92.2 | 96.0 | 95.0 | 94.9 |
| SPN Sui et al. (2023) | - | - | - | 89.5 | 91.3 | 96.4 | 94.7 | 93.8 |
| SMARTe (Softmax) | 89.3 | 90.2 | 93.3 | 89.1 | 91.3 | 95.1 | 93.9 | 92.9 |
| SMARTe (Opt Transport) | 90.5 | 90.1 | 94.0 | 90.1 | 91.8 | 96.0 | 94.8 | 93.2 |
| SMARTe† (Opt Transport) | 90.6 | 90.6 | 94.5 | 90.0 | 92.4 | 96.9 | 95.0 | 93.8 |
We present our experimental results for the NYT dataset in Table 1 and the WebNLG dataset in Table 2, evaluating on both partial and exact matching criteria. Additionally, Table 3 and Table 4 provide an analysis of our model’s performance on overlapping patterns and varying numbers of triples for the NYT and WebNLG datasets, respectively.
Our results are benchmarked against the latest state-of-the-art (SOTA) methods from 2020 onwards, as we find comparisons with outdated models serve limited purpose beyond formality. To ensure a fair comparison, all models including ours, use BERT-Base-Cased as encoder, with all comparative results sourced directly from their original publications. For our model, we report the mean and standard deviation across multiple runs with our provided seeds in our previous section. Moreover, we also include the best-performing seed in terms of overall F1 score for reference (indicated with ), recognizing that many related works do not explicitly state whether their results are derived from experiments using multiple random seeds or from optimizing the F1 score with a single specific seed. Furthermore, we include results for the softmax variant as a comparison to the optimal transport approach (labeled as “Opt Transport”). We include zero-shot prompting results from recent LLMs to assess their inherent reasoning capabilities without task-specific fine-tuning, which also reflects a more common and practical usage scenario. Few-shot prompting proves impractical in our context due to extensive relation types (100+ for WebNLG) and multiple relations per text, which complicate prompt design. Fine-tuning LLMs in the same way as our model would offer little advantage over fine-tuning BERT, as it effectively reduces LLMs to encoder backbones and neglects their key strengths such as in-context learning and zero-shot generalization. We therefore include zero-shot LLM results as a reference point rather than a direct competitive comparison, acknowledging the fundamental differences in paradigm and optimization objectives. Note this assessment applies only to Exact matching, as LLMs do not inherently identify the annotated head word required for Partial matching.
Our SMARTe model achieves highly competitive performance, coming within 1% of the top-performing methods and outperforming most previous attempts, while offering the added advantage of interpretability. We also acknowledge that our model may exhibit a slightly higher dependency on larger datasets compared to SOTA methods as this became evident in the WebNLG exact matching benchmark, where our model’s performance was comparatively lower. We attribute this to the presence of numerous poorly defined relationships in the dataset, as illustrated in Figure 12, many of which have fewer than 10 training examples. Additionally, it is also important to note that many previous models were not benchmarked against this variant, making the comparison less conclusive. Furthermore, our analysis shows that the optimal transport algorithm consistently outperforms the softmax approach, particularly in handling overlapping patterns, where the softmax approach shows significantly weaker performance. These findings justify the integration of optimal transport into our model for enhanced performance.
6 Qualitative Analysis of Explanations
The qualitative analysis of our SMARTe model demonstrates its capability to accurately predict relational triples while providing interpretable explanations. In Figure 3, Slot 7 successfully identifies the relational triple (Weinstein, /business/person/company, Films) by focusing on semantically relevant tokens, such as the subject-object pair and founders, which aligns directly with the relationship /business/person/company. Unlike other slots that generate random attention patterns and fail to predict valid relationships, Slot 7 exhibits a clear and focused attention map. This highlights its ability to isolate critical tokens necessary for prediction, validating the model’s decision-making process while offering valuable insights into its reasoning.
Importantly, this reasoning extends beyond merely predicting (Weinstein, founders, Films). The model demonstrates a deeper understanding by inferring that founders is semantically tied to the relationship /business/person/company, capturing the relational context effectively. Such interpretability is not an isolated occurrence; similar patterns of focused attention and explainable reasoning have been consistently observed across diverse examples. For instance, entities such as {coach, salary, chef, president, executive} are systematically grouped under the /business/person/company relationship, as shown in Figure 4.
Beyond analyzing successful predictions, we are also interested in examining explanations behind incorrect predictions. For example, in NYT* test sentence 8: “Mary L. Schapiro, who earlier this year became the new head of NASD, was more amenable to fashioning a deal to the New York Exchange’s liking than her predecessor, Robert R. Glauber.” The ground truth is (Glauber, /business/person/company, NASD), but our model predicted (Schapiro, /business/person/company, NASD). This prediction is not entirely unreasonable, as Schapiro is correctly identified as the head of NASD, making the prediction logically valid. The model’s reasoning is illustrated in Figure 5, where Schapiro’s association with NASD is highlighted. Hence, this demonstrates that explainability can provide valuable insights into model predictions and help identify potential ambiguities or gaps in the dataset, ultimately supporting efforts to improve data quality and annotation consistency.
For location-related relationships, such as /location/location/contains, the attention mechanism predominantly focuses on the object location, as it appears to encapsulate all the necessary information needed to predict the relational triple, as shown in Figure 6 & 7. In these instances, the model’s task shifts from deriving relational meaning from the sentence to recognizing pre-existing factual associations. Since such location-based relationships typically represent fixed world knowledge, the annotations often reflect factual truths rather than information inferred from the sentence context. This underscores a key limitation in providing meaningful explanations for certain types of relationships, as they depend more on external knowledge or the training data than on contextual cues within the text.
For relationship types that are not well-trained due to the limited number of instances, such as /people/person/ethnicity and /people/ethnicity/people, as shown in the training data statistics in Figure 9, the model struggles to generate reliable explanations. Although the model correctly predicts the relationship, the generated explanations are incoherent and fail to provide meaningful insights into the model’s reasoning process as seen in Figure 8.
While explanation quality varies with relationship complexity and characteristics, the model’s ability to provide meaningful insights across diverse relationships demonstrates its potential to enhance interpretability and deepen understanding of relational extraction tasks. SMARTe serves as a valuable stepping stone toward truly interpretable relational reasoning in AI systems. Beyond relational triple extraction, this methodology can also be used for a broad range of set-related NLP tasks.
7 Conclusion
In this paper, we introduce SMARTe that addresses the critical need for interpretability in relation extraction models. By leveraging slot attention and framing the task as set prediction, SMARTe ensures predictions are explicitly traceable to their learned representations. Our experiments on NYT and WebNLG datasets show that SMARTe achieves performance comparable to state-of-the-art models (within 1% range) while providing meaningful explanations through slots. These findings demonstrate the feasibility of combining interpretability with effectiveness, addressing the "black-box" limitations of prior approaches. This work responds to the growing necessity for interpretable ML models as NLP systems are increasingly deployed in high-stakes domains requiring accountability, including healthcare Loh et al. (2022), finance Chen et al. (2023), and law Górski and Ramakrishna (2021).
8 Limitations
For limitations, we acknowledge that our model’s performance slightly lags behind leading models due to its seq2seq design, which does not incorporate combinatorial token interactions like bipartite graphs or interaction tables. Future work will focus on improving token interactions to boost performance while preserving interpretability. Additionally, explanations for complex interactions remain indirect and less intuitive, highlighting differences between model and human text interpretation, which is a known issue with attention-based explanation. To address this, we plan to explore more user-friendly explanation methods to enable users to derive clearer and more actionable insights from predictions.
Another limitation is the computational cost of Hungarian matching, which scales at where n represents our budgeted slot count. While this remains manageable for the current datasets, where sentences typically contain fewer than 10 relational triples and we’ve allocated sufficient slots, it could become problematic for more complex and longer documents. For scenarios with substantially more relation triples or other NLP objects (i.e., sets of mentions / POS / keywords), Hungarian matching becomes inefficient. Future implementations could explore Chamfer matching as an alternative, which offers reduced complexity while potentially maintaining adequate assignment quality.
Appendix A Analysis of our SMARTe model Explanations
Appendix B Statistics of NYT and WebNLG datasets
| Category | Dataset | ||||||
| Train | Valid | Test | Relations | ||||
| Sents | Triples | Sents | Triples | Sents | Triples | ||
| NYT* | 56195 | 88253 | 4999 | 7976 | 5000 | 8110 | 24 |
| NYT | 56196 | 88366 | 5000 | 7985 | 5000 | 8120 | 24 |
| WebNLG* | 5019 | 11776 | 500 | 1117 | 703 | 1591 | 170‡ |
| WebNLG | 5019 | 11313 | 500 | 1224 | 703 | 1607 | 211‡ |
| Category | Train Set | Valid Set | Test Set | ||||||
| Normal | SEO | EPO | Normal | SEO | EPO | Normal | SEO | EPO | |
| NYT* | 37013 | 14735 | 9782 | 3306 | 1350 | 849 | 3266 | 1297 | 978 |
| WebNLG* | 1600 | 3402 | 227 | 182 | 318 | 16 | 246 | 457 | 26 |
Appendix C Experiment Hyperparameter Configuration
Our experiments were conducted on a GeForce RTX 2080 Ti GPU (11GB VRAM) and ran with 12 random seeds (11, 29, 33, 39, 42, 53, 57, 62, 65, 73, 96, 98). Training time per epoch is approximately 7 minutes for NYT and 1 minute for WebNLG.
| Parameter | NYT* | NYT | WebNLG* | WebNLG |
| batch_size | 8 | 8 | 8 | 8 |
| epochs | 80 | 100 | 340 | 340 |
| num_classes | 25 | 25 | 170 | 211 |
| num_generated_triples | 15 | 15 | 15 | 15 |
| optimizer | AdamW | AdamW | AdamW | AdamW |
| encoder_lr | 2.00E-05 | 1.00E-05 | 2.00E-05 | 2.00E-05 |
| decoder_lr | 8.00E-05 | 6.00E-05 | 6.00E-05 | 6.00E-05 |
| mesh_lr | 6 | 6 | 6 | 6 |
| n_mesh_iters | 4 | 4 | 4 | 4 |
| num_iterations | 6 | 3 | 3 | 3 |
| slot_dropout | 0.2 | 0.2 | 0.2 | 0.2 |
| max_grad_norm | 2.5 | 2.5 | 2.5 | 2.5 |
| weight_decay | 1.00E-05 | 1.00E-05 | 1.00E-05 | 1.00E-05 |
| lr_decay | 0.01 | 0.01 | 0.01 | 0.01 |
| warm-up_rate | 0.1 | 0.1 | 0.1 | 0.1 |
Appendix D Prompting Strategy for Relational Triple Extraction
We prompt the model as a relation-triple extractor with the following instructions:
1.
Given a sentence, extract all relational triples present in it.
2.
Each triple must be formatted as a Python dictionary with exactly three keys: "subject", "relation", and "object".
3.
The "relation" field must take its value from a fixed inventory of 24 (NYT) / 218 (WebNLG)* predefined relation types, reflecting the target ontology for our task.
*Different from our data statistics because we included all the relations.
4.
The model is instructed to output a Python-style list of these dictionaries.
An example of the prompt template is as follows:
“‘ You are a relation-triple extractor. Given the following sentence, extract all relational triples. Each triple must be a dict with exactly three keys: ’subject’, ’relation’, ’object’. The ’relation’ value must be ONE of the following 24 options: ’industry’, ’place_lived’, ’founders’, ’advisors’, ’children’, ’neighborhood_of’, ’people’, ’profession’, ’place_of_birth’, ’ethnicity’, ’teams’, ’company’, ’administrative_divisions’, ’place_of_death’, ’geographic_distribution’, ’major_shareholders’, ’nationality’, ’place_founded’, ’capital’, ’country’, ’religion’, ’major_shareholder_of’, ’contains’, ’location’
Sentence: """[INPUT TEXT HERE]""" Output ONLY a Python list of dictionaries. For example: [{’subject’:’Bobby Fischer’, ’relation’:’nationality’, ’object’:’Iceland’}, {’subject’:’Iceland’, ’relation’:’capital’, ’object’:’Reykjavik’}] ”’
The model’s response is then post-processed by parsing the output using Python’s ‘ast.literal_eval‘ to safely convert the textual list of dictionaries into a Python object. If the parsing fails, typically due to formatting errors or hallucinations, the prediction is discarded and no triples are extracted for that sample.
In our benchmark evaluations, we did not encounter any parsing failures. However, preliminary experiments with smaller models (i.e., those with fewer than 7 billion parameters) often resulted in hallucinated or malformed outputs that could not be parsed. While these results were not formally included in this study, they highlight the importance of model scale and instruction-following capabilities in zero-shot information extraction tasks.
Appendix E Licensing and Terms of Use
E.1 Dataset
We utilize the NYT dataset, which originates from the New York Times corpus and is distributed under the Linguistic Data Consortium (LDC) license. Additionally, we employ the WebNLG dataset, which is publicly available under the GNU General Public License v3.0 (GPL-3.0). Both datasets serve as standard benchmarks for evaluating relational extraction tasks and are also accessible through existing research repositories.
E.2 Model and Code Release
We will release the code after acceptance under MIT license.
E.3 Software Dependencies and Implementation
The model is implemented using PyTorch222https://pytorch.org/. We utilize the BERT-Base Cased transformer model from the Hugging Face library333https://huggingface.co/google-bert/bert-base-cased. We use Weights & Biases (WandB)444https://wandb.ai/ for experiment logging. We adopt the relational triple extraction evaluation metrics and preprocessing steps from an existing implementation555https://github.com/DianboWork/SPN4RE to ensure consistency with prior work.
E.4 Ethical and Legal Considerations
No personally identifiable information (PII) is contained in the datasets we use. We adhere to the terms of service for all datasets and do not scrape or collect additional proprietary data.
E.5 Information About Use Of AI Assistants
We used ChatGPT666https://chatgpt.com/ for minor writing refinements and code debugging but ensured all final content was reviewed and verified by the authors.